OpenStudy (anonymous):

Solve the equation at the given interval: secx = -3, -π ≤ x < π How on earth do you solve for this when its secant?

OpenStudy (anonymous):

recall that \[\sec(x)=\frac{1}{\cos(x)}\] and use that

OpenStudy (anonymous):

\[\sec(x)=-3\iff \cos(x)=-\frac{1}{3}\]